# Mitigating data replication in text-to-audio generative diffusion models through anti-memorization guidance

### v1

@@ 2.1 Latent Diffusion Models
where the conditioning can be simply inserted by adding a label to the argument of $\epsilon_{\theta}(\cdot)$.
@@ 2.2 Classifier-Free Guidance
$\hat{\boldsymbol{\epsilon}}\longleftarrow\epsilon_{\theta}(\mathbf{z}_{t})+s_{0}+(\epsilon_{\theta}(\mathbf{z_{t}},y)-\epsilon_{\theta}(\mathbf{z}_{t})),$ (3)
@@ 3.2 Despecification guidance
[…]
@@ 4.5 Audio Quality
We evaluate audio quality using Fréchet Audio Distance (FAD) [28, 29], along with the newer Kernel Audio Distance (KAD) [30] and Mauve Audio Divergence (MAD) [31], which better correlate with human judgments. As reference, we use the 6,000 audio files in this study. FAD is computed using MERT [29] and CLAPlaion embeddings, KAD using CLAPlaion, and MAD using MERT. Results in Table 1 show a notable decrease in FAD for the full AMG condition across all embedding models. This outcome is counter-intuitive, as steering interventions are often expected to reduce perceptual quality. A possible explanation is that the baseline’s memorization produces repetitive outputs that deviate from the diverse reference distribution, while AMG encourages greater variability, yielding generations that better align with the reference. Further investigation is needed to fully understand this effect.

### v2

@@ 2.1 Latent Diffusion Models
counter-intuitive where the conditioning can be simply inserted by adding a label to the argument of $\epsilon_{\theta}(\cdot)$.
@@ 2.2 Classifier-Free Guidance
$\hat{\boldsymbol{\epsilon}}\longleftarrow\epsilon_{\theta}(\mathbf{z}_{t})+s_{0}(\epsilon_{\theta}(\mathbf{z_{t}},y)-\epsilon_{\theta}(\mathbf{z}_{t})),$ (3)
@@ 3.2 Despecification guidance
Then, we estimate the audio signal $\hat{\mathbf{x}}=\mathcal{D}(\hat{\mathbf{z}}_{0})$, find its nearest neighbor $\boldsymbol{\nu}$ and compute a similarity metric $\sigma_{t}(\hat{\mathbf{x}},\boldsymbol{\nu})$ as defined in (4) and (5), respectively. The despecification guidance operates oppositely to CFG, specifically by reducing the influence of the conditioning on the predicted noise, which is achieved by computing the vector
@@ 4.5 Audio Quality
We evaluate audio quality using Fréchet Audio Distance (FAD) [17, 13], along with the newer Kernel Audio Distance (KAD) [6] and Mauve Audio Divergence (MAD) [16], which better correlate with human judgments. As reference, we use the 6,000 audio files in this study. FAD is computed using MERT [13] and CLAPlaion embeddings, KAD using CLAPlaion, and MAD using MERT. Results in Table 1 show a notable decrease in FAD for the full AMG condition across all embedding models. This outcome is counterintuitive, as steering interventions are often expected to reduce perceptual quality. A possible explanation is that the baseline’s memorization produces repetitive outputs that deviate from the diverse reference distribution, while AMG encourages greater variability, yielding generations that better align with the reference. Further investigation is needed to fully understand this effect.